# A Survey of Controllable Learning: Methods and Applications in Information Retrieval

### v2

@@ 1 Introduction
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=base east, child anchor=base west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=13em,fill=fill_0,draw=draw_0,font=,, where level=1text width=9.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=10.4em,fill=fill_2,draw=draw_2,font=,, where level=3text width=10.4em,fill=fill_3,draw=draw_3,font=,, where level=4font=,, [ Controllable Learning (CL)
, minimum height=3em [ Formulation of CL
@@
(Sec. 2) [ Define a task requirement triplet, $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. ,leaf, text width=31.8em ] ] ] [ Taxonomy of CL
@@
(Sec. 3.1.1) [ Taking several objectives (metrics) as control targets, these work includes ComiRec [1], UCRS [2], CMR [3], PadiRec [4], RecLM-gen [5] , leaf, text width=31.8em ] ] [ User Portrait Control
(Sec. 3.1.2) [ Taking historical behavior as control targets , these work includes LACE [6], LFRQE [7], Grapevine [8], TEARS [9], CMBR [10], UCR [11] , leaf, text width=31.8em ] ] [ Scenario Adaptation Control
(Sec. 3.1.3) [ Taking environment adaptation as target, these work includes HyperBndit[12], Hamur [13], HyperBandit [12], PEPNet [14] , leaf, text width=31.8em ] ] ] [ Who controls
@@
(Sec. 3.2.1) [ Prior profile modification which empowers users to actively shape their recommendation experience: UCRS [2], IFRQE [7], LACE [6], Supervised $\beta$-VAE [15], LP [16], TEARS [9], LangPTune [17],UCRS [2] , leaf, text width=31.8em ] ] [ Platform-mediated control
(Sec. 3.2.2) [ Impose algorithmic adjustments and policy-based constraints on the recommendation process: ComiRec [1], CMR [3], CCDF [18] and  [19], Padirec [4], SAMD [20], HyperBandit [12], DTRN [21] , leaf, text width=31.8em ] ] ] ] [ How control is
@@
(Sec. 4.1.1) [ Rule-based post-processing techniques: [22], [23], [24], [25], [26] and [27]. , leaf, text width=31.8em ] ] [ Pareto optimization
(Sec. 4.1.2) [ Achieving controllability by using Pareto Optimization: PHN-HVI [28], PHNs [29],  [30],  [31],  [32],  [33],  [34] and  [35]. , leaf, text width=31.8em ] ] [ Hypernetwork
(Sec. 4.1.3) [ Hypernetworks based methods: Hamur [13], CMR [3], Hyperprompt [36], HyperBandit [12], PadiRec [4], [37], [28], and [29] , leaf, text width=31.8em ] ] ] [ Where to control
(Sec. 4.2) [ Pre-processing methods
(Sec. 4.2.1) [ Including data manipulation methods and feature engineering methods: UCRS [2], LACE [6],  [7] and MetaLens [38]. , leaf, text width=31.8em ] ] [ In-processing methods
(Sec. 4.2.2) [ Including regularization and constrained optimization methods:CMR [3], CCDF [18], HyperBandit [12] , leaf, text width=31.8em ] ] [ Post-processing methods
(Sec. 4.2.3) [ Including reranking and result diversification methods: ComiRec [1] and MMR [39] , leaf, text width=31.8em ] ] ] ] [ Evaluations for CL
in IR (Sec. 5) [ Metrics (Sec. 5.1), minimum height=3em[ Single-objective metrics
(Sec. 5.1.1) [ NDCG [40], Recall,  Precision,  Hit Rate,  $\alpha$-NDCG [41], ERR-IA [42], Coverage, Iso-Index. , leaf, text width=31.8em ] ] [ Multi-objective metrics
(Sec. 5.1.2) [ Typical Multi-objective metrics include 4 main groups [43]. However, only the Hypervolume indicator [44] is most suitable in the current discussion context. , leaf, text width=31.8em ] ] ] [ Datasets (Sec. 5.2), minimum height=3em[ Amazon [45, 46], Ali_Display_Ad_Click [47], UserBehavior [1], MovieLens, MS MARCO [48] , leaf, text width=31.8em ] ] ] [ Challenges in CL
@@
(Sec. 6.1)[ Pursuing controllability often leads to a trade-off, potentially compromising performance or other user-centric optimization metrics and adversely impacting accuracy or user experience. ,leaf, text width=31.8em ] ] [ Absence of evaluation
(Sec. 6.2)[ The assortment of perspectives on controllability and the consequent need for tailored evaluation metrics prevents direct methodological comparisons and hinder the progression of the field. , leaf, text width=31.8em ] ] [ Descriptions setting
challenges (Sec. 6.3)[ The task description serves as the instructions given by humans to the learner. A crucial issue is how to set the task target and transform it into a human understandable and precise description. ,leaf, text width=31.8em ] ] [ Online challenges
(Sec. 6.4)[ Scalability in realworld IR systems, particularly those dealing with streaming data and requiring continuous learning, is a formidable challenge. , leaf, text width=31.8em ] ] ] [ Future directions of
@@
Figure 1: Overview of the survey of controllable learning (CL). This survey includes formulation of CL, taxonomy of CL, implementation methods of CL, how to evaluate the effectiveness of CL in information retrieval (IR) applications, and challenges of applying CL in IR.
@@
Trustworthiness in machine learning is becoming increasingly important, with considerations such as controllability, fairness, privacy, and interpretability gaining prominence [49, 50, 51]. In November 2023, the Bletchley Declaration [52] , adopted at the first Global Artificial Intelligence Safety Summit, emphasized that:“AI should be designed, developed, deployed, and used, in a manner that is safe, in such a way as to be human-centric, trustworthy and responsible…Substantial risks may arise from potential intentional misuse or unintended issues of control relating to alignment with human intent.” In October 2023, the Cyberspace Administration of China released the Global AI Governance Initiative [53], which emphasized the need to: “…make AI technologies more secure, reliable, controllable, and equitable, …ensure that AI always remains under human control.” In this paper, we explore the topic of controllable AI, with a particular focus on the controllability of machine learning methods and their practical applications. Norbert Wiener, the originator of cybernetics, foresaw the ethical challenge of learning machine over sixty years ago when he stated [54]As machines learn they may develop unforeseen strategies at rates that baffle their programmers.” Additionally, Wiener explained the significance of controllability in learning machines [54]: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot efficiently interfere once we have started it, because the action is so fast and irrevocable that we have not the data to intervene before the actions complete, then we had better be quite sure that the purpose put into the machine is the purpose which we really desire and not merely a colorful imitation of it, etc.”, emphasizing two fundamental aspects of AI controllability: (1) A controllable learning machine needs to ensure that its predictions/decisions meet the target of the AI user; (2) Intervening in a learning machine after its deployment is a crucial but challenging task.
@@
Controllability has been extensively discussed in the field of generative machine learning models, also known as controllable generation. For instance, in text generation [55, 56, 57] or visual generation [58, 59, 60], the AI models take a given task description (often referred to as a “prompt”) as input and generates content that aligns with that description. However, in the realm of discriminative machine learning models and applications, there lacks a unified definition and in-depth discussion of controllability. The current popular prompt-based controllable generation methods can be seen as a form of pre-processing, where, under fixed model parameters, changing input features allows the model to meet the target of the AI user. Existing theoretical results suggest that controlling model parameters may lead to better convergence than controlling model inputs when given task requirements [37], revealing that achieving controllability in machine learning involves various methods beyond prompt-based approaches. Therefore, controllable learning (CL) is a crucial category worthy of comprehensive review and exploration in trustworthy machine learning, ensuring that learning models meet user expectations and adhere to ethical standards. Meanwhile, in the context of learning-based Information Retrieval (IR) applications, users actively search for information or passively receive recommendations, both aimed at fulfilling their information-seeking needs [61, 62]. In these applications, controllable learning (CL) enables models to adapt to to diverse task description (e.g, users’ information needs) at test time without extensive retraining, delivering personalized and relevant results. Given the above landscape, this survey attempts to provide a formal definition of controllable learning, classify and discuss its application paradigms in information retrieval. In summary, this survey focuses on the topic of controllable learning, which aims to find a learner capable of adapting to different task requirements without the need for retraining, thereby meeting the desired task targets of the AI users.
@@
In the context of MaaS, the demand for controllability of ML models in information retrieval applications becomes more pronounced. This is attributed to the inherent complexity of information needs in IR tasks, where precise and tailored results are critical. In traditional settings, such as keyword-based search engines, the retrieval process is relatively straightforward, relying on predefined rules and algorithms. However, with the advent of MaaS and the utilization of large language models, the landscape has evolved. Now, users expect finer-grained and context-aware responses based on their information needs. This heightened expectation emphasizes the importance of controllability of ML models. Specifically, there is a need for mechanisms that enable individuals or platforms to easily express and convey their specific requirements for IR tasks. For example, in an e-commerce setting, a user searching for “black T-shirt” may expect different results based on varying weights of evaluation metrics (such as favoring diversity to simultaneously expose clothing that pairs well with black T-shirts), or setting specific filters to exclude irrelevant content (such as not exposing previously purchased T-shirts). Achieving this level of controllability necessitates sophisticated ML models capable of dynamically adjusting model parameters or outputs based on task requirements, without extensive retraining. Therefore, in the realm of IR, MaaS not only provides unprecedented access to powerful ML models but also underscores the significance of controllability and customization in addressing the diverse needs of users across different domains and applications.
The current landscape of research and application deployment in the aforementioned areas motivates us to write a comprehensive survey on controllable learning. As shown in Figure 1, this survey will encompass various aspects including the formal definition of controllable learning, methods and applications in information retrieval, evaluations, challenges, and future directions.
  • Formal definition of controllable learning. The survey will begin by providing a clear and concise formal definition of controllable learning.
  • Applications of controllable learning in IR. We will then delve into the applications of controllable learning in the field of IR, exploring who controls, what is controllable, and how control is implemented in IR systems.
  • Future directions and challenges. Finally, the survey will discuss the future directions and potential challenges in the field of controllable learning.
Overall, the survey on controllable learning will serve as a valuable resource for students, researchers, practitioners, and policymakers interested in understanding the principles, applications, and future prospects of this important area of trustworthy machine learning.
@@ 1.1 Comparison with Existing Surveys
The survey on trustworthy recommender systems [66] also talks about controllability, which divides it into two distinct categories: explicit controllability and implicit controllability. The former allows users explicitly edit or update the user preferences, indicating the users are exactly aware of their favors. The latter means that users could indirectly fine-tune their preferences when dynamically interacting with the recommender systems through re-ranking, modifying historical information, and so on. Similarly, a survey on user control in recommender systems [67] also split the controllability into two parts, where one is control during the preference elicitation phase and the other is control in the context of recommendation results. These two parts is corresponding to explicit controllability and implicit controllability above.
@@
The aforementioned surveys primarily focus on recommender systems where users either explicitly adjust their preferences or implicitly allow the system to infer preferences from their behavior. However, they lack a summary of how controllability can be formally defined in machine learning. Moreover, there may be other possibilities for controllability in a broader sense, such as platform-side controllability or controllability with multiple objectives, which are not thoroughly defined and categorized. These surveys are also relatively early, and many new controllable techniques (such as hypernetworks [68], large language models [63], etc.) have emerged since then. Considering the limitations of existing surveys, in this survey, we provide a formal definition of controllable learning, and discuss controllable learning and its applications in information retrieval.
2 Foundations of Controllable Learning

### v3

@@ 1 Introduction
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=base east, child anchor=base west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=13em,fill=fill_0,draw=draw_0,font=,, where level=1text width=9.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=10.4em,fill=fill_2,draw=draw_2,font=,, where level=3text width=10.4em,fill=fill_3,draw=draw_3,font=,, where level=4font=,, [ Controllable Learning
(CL) , minimum height=3em [ Formulation of CL
@@
(Sec. 2) [ Define a task requirement triplet, $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. ,leaf, text width=31em ] ] ] [ Taxonomy of CL
@@
(Sec. 3.2.1) [ Taking several objectives (metrics) as control targets, these work includes ComiRec [1], UCRS [2], CMR [3], PadiRec [4], RecLM-gen [5] , leaf, text width=31em ] ] [ User Portrait Control
(Sec. 3.2.2) [ Taking historical behavior as control targets , these work includes LACE [6], LFRQE [7], Grapevine [8], TEARS [9], CMBR [10], UCR [11] , leaf, text width=31em ] ] [ Scenario Adaptation
Control (Sec. 3.2.3) [ Taking environment adaptation as target, these work includes HyperBndit[12], Hamur [13], HyperBandit [12], PEPNet [14] , leaf, text width=31em ] ] ] [ Who controls
@@
(Sec. 3.3.1) [ Prior profile modification which empowers users to actively shape their recommendation experience: UCRS [2], IFRQE [7], LACE [6], Supervised $\beta$-VAE [15], LP [16], TEARS [9], LangPTune [17],UCRS [2], GOMMIR [18], Promptriever [19], InstructAgent [20] , leaf, text width=31em ] ] [ Platform-mediated control
(Sec. 3.3.2) [ Impose algorithmic adjustments and policy-based constraints on the recommendation process: ComiRec [1], CMR [3], CCDF [21] and  [22], Padirec [4], SAMD [23], HyperBandit [12], DTRN [24] , leaf, text width=31em ] ] ] ] [ How control is
@@
(Sec. 4.1.1) [ Rule-based post-processing techniques: [25], [26], [27], [28], [29] and [30]. , leaf, text width=31em ] ] [ Pareto optimization
(Sec. 4.1.2) [ Achieving controllability by using Pareto Optimization: PHN-HVI [31], PHNs [32],  [33],  [34],  [35],  [36],  [37] and  [38]. , leaf, text width=31em ] ] [ Hypernetwork
(Sec. 4.1.3) [ Hypernetworks based methods: Hamur [13], CMR [3], Hyperprompt [39], HyperBandit [12], PadiRec [4], Hypencoder [40], [41], [31], and [32] , leaf, text width=31em ] ] ] [ Where to control
(Sec. 4.2) [ Pre-processing
methods(Sec. 4.2.1) [ Including data manipulation methods and feature engineering methods: UCRS [2], LACE [6],  [7] and MetaLens [42]. TART [43], Instructor [44], FollowIR [45], InstructIR [46] , leaf, text width=31em ] ] [ In-processing
methods(Sec. 4.2.2) [ Including regularization and constrained optimization methods:CMR [3], CCDF [21], HyperBandit [12] , leaf, text width=31em ] ] [ Post-processing
methods(Sec. 4.2.3) [ Including reranking and result diversification methods: ComiRec [1] and MMR [47] , leaf, text width=31em ] ] ] ] [ Evaluations for CL
in IR (Sec. 5) [ Metrics (Sec. 5.1), minimum height=3em[ Single-objective
metrics(Sec. 5.1.1) [ NDCG [48], Recall,  Precision,  Hit Rate,  $\alpha$-NDCG [49], ERR-IA [50], Coverage, Iso-Index. , leaf, text width=31em ] ] [ Multi-objective
metrics(Sec. 5.1.2) [ Typical Multi-objective metrics include 4 main groups [51]. However, only the Hypervolume indicator [52] is most suitable in the current discussion context. , leaf, text width=31em ] ] ] [ Datasets (Sec. 5.2), minimum height=3em[ Amazon [53, 54], Ali_Display_Ad_Click [55], UserBehavior [1], MovieLens, MS MARCO [56] , leaf, text width=31em ] ] ] [ Challenges in CL
@@
(Sec. 6.1)[ Pursuing controllability often leads to a trade-off, potentially compromising performance or other user-centric optimization metrics and adversely impacting accuracy or user experience. ,leaf, text width=31em ] ] [ Absence of evaluation
(Sec. 6.2)[ The assortment of perspectives on controllability and the consequent need for tailored evaluation metrics prevents direct methodological comparisons and hinder the progression of the field. , leaf, text width=31em ] ] [ Descriptions setting
challenges (Sec. 6.3)[ The task description serves as the instructions given by humans to the learner. A crucial issue is how to set the task target and transform it into a human understandable and precise description. ,leaf, text width=31em ] ] [ Online challenges
(Sec. 6.4)[ Scalability in realworld IR systems, particularly those dealing with streaming data and requiring continuous learning, is a formidable challenge. , leaf, text width=31em ] ] ] [ Future directions of
@@
Figure 1: Overview of the survey of CL. This survey includes formulation of CL, taxonomy of CL, implementation methods of CL, how to evaluate the effectiveness of CL in information retrieval (IR) applications, and challenges of applying CL in IR.
@@
Trustworthiness in machine learning is becoming increasingly important, with considerations such as controllability, fairness, privacy, and interpretability gaining prominence [57, 58, 59, 60]. In November 2023, the Bletchley Declaration , adopted at the first Global Artificial Intelligence Safety Summit, emphasized that:“AI should be designed, developed, deployed, and used, in a manner that is safe, in such a way as to be human-centric, trustworthy and responsible…Substantial risks may arise from potential intentional misuse or unintended issues of control relating to alignment with human intent.” In October 2023, the Cyberspace Administration of China released the Global AI Governance Initiative, which emphasized the need to: “…make AI technologies more secure, reliable, controllable, and equitable, …ensure that AI always remains under human control.” In this paper, we explore the topic of controllable AI, with a particular focus on the controllability of machine learning methods and their practical applications. Norbert Wiener, the originator of cybernetics, foresaw the ethical challenge of learning machine over sixty years ago when he stated [61]As machines learn they may develop unforeseen strategies at rates that baffle their programmers.” Additionally, Wiener explained the significance of controllability in learning machines [61]: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot efficiently interfere once we have started it, because the action is so fast and irrevocable that we have not the data to intervene before the actions complete, then we had better be quite sure that the purpose put into the machine is the purpose which we really desire and not merely a colorful imitation of it, etc.”, emphasizing two fundamental aspects of AI controllability: (1) Controllable learning (CL) machine needs to ensure that its predictions/decisions meet the target of the AI user; (2) Intervening in a learning machine after its deployment is a crucial but challenging task.
Under this insight and context, we present an analysis of the feasibility and necessity of CL from three perspectives: the technical foundations of controllability, the practical value of CL in information retrieval, and the landscape of commercial deployment.
1.1 Technical Foundations of Controllable Learning (CL)
@@
Controllability has been extensively discussed in the field of generative machine learning models, also known as controllable generation. For instance, in text generation [62, 63, 64] or visual generation [65, 66, 67], the AI models take a given task description (often referred to as a “prompt”) as input and generates content that aligns with that description. However, in the realm of discriminative machine learning models and applications, there lacks a unified definition and in-depth discussion of controllability. The current popular prompt-based controllable generation methods can be seen as a form of pre-processing, where, under fixed model parameters, changing input features allows the model to meet the target of the AI user. Existing theoretical results suggest that controlling model parameters may lead to better convergence than controlling model inputs when given task requirements [41], revealing that achieving controllability in machine learning involves various methods beyond prompt-based approaches. Therefore, controllable learning (CL) is a crucial category worthy of comprehensive review and exploration in trustworthy machine learning, ensuring that learning models meet user expectations and adhere to ethical standards.
1.2 Practical Value of CL in Information Retrieval
Building on the technical foundations of CL, it is important to examine how these principles translate into practical benefits within real-world machine learning applications. One such domain is Information Retrieval (IR). Specifically, in the context of learning-based Information Retrieval (IR) applications, users actively search for information or passively receive recommendations, both aimed at fulfilling their information-seeking needs [68, 69, 70, 71, 72]. In these applications, CL enables models to adapt to to diverse task description (e.g, users’ information needs) at test time without extensive retraining, delivering personalized and relevant results. Given the above landscape, this survey attempts to provide a formal definition of CL, classify and discuss its application paradigms in information retrieval. In summary, this survey focuses on the topic of CL, which aims to find a learner capable of adapting to different task requirements without the need for retraining, thereby meeting the desired task targets of the AI users.
1.3 Landscape of Commercial Deployment
@@
In the context of MaaS, the demand for controllability of ML models in information retrieval applications becomes more pronounced. This is attributed to the inherent complexity of information needs in IR tasks, where precise and tailored results are critical. In traditional settings, such as keyword-based search engines, the retrieval process is relatively straightforward, relying on predefined rules and algorithms. However, with the advent of MaaS and the utilization of large language models, the landscape has evolved. Now, users expect finer-grained and context-aware responses based on their information needs [76, 77]. This heightened expectation emphasizes the importance of controllability of ML models. Specifically, there is a need for mechanisms that enable individuals or platforms to easily express and convey their specific requirements for IR tasks. For example, in an e-commerce setting, a user searching for “black T-shirt” may expect different results based on varying weights of evaluation metrics (such as favoring diversity to simultaneously expose clothing that pairs well with black T-shirts), or setting specific filters to exclude irrelevant content (such as not exposing previously purchased T-shirts). Achieving this level of controllability necessitates sophisticated ML models capable of dynamically adjusting model parameters or outputs based on task requirements, without extensive retraining. Therefore, in the realm of IR, MaaS not only provides unprecedented access to powerful ML models but also underscores the significance of controllability and customization in addressing the diverse needs of users across different domains and applications.
@@ 1.4 Comparison with Existing Surveys
This section aim to state the distinctions between this survey and existing surveys. To the best of our knowledge, there is a notable lack of comprehensive surveys specifically dedicated to controllable learning in information retrieval. To date, only one survey from 2017 has touched upon the concept of “user control” as a keyword [78]. Therefore, we compare our work with surveys in closely related areas, including Trustworthy Information Retrieval [79], and Explainable Information Retrieval [80, 81]
To begin with, the survey on user control [78] focuses on the instantaneous nature of user interests. It highlights the need for mechanisms that allow users to take control of recommendations when the system’s assumptions about their preferences are inaccurate or outdated. However, the survey exhibits significant limitations. It considers controllability exclusively from the user’s perspective, neglecting the role of platform-side control. Moreover, it merely summarizes existing work in terms of interaction forms, lacking rigorous definitions and in-depth technical analysis. Similarly, the survey on trustworthy information retrieval [79] primarily focuses on the security aspects of privacy preservation in recommender systems. Nevertheless, it also addresses controllability, categorizing it into two distinct types: explicit controllability and implicit controllability. The former allows users explicitly edit or update the user preferences. The latter means that users could indirectly fine-tune their preferences when dynamically interacting with the recommender systems through re-ranking, modifying historical information, and so on. However, this survey lacks a systematic summary of the definition of controllability in machine learning and the methods by which controllability can be achieved. Moreover, there may be other possibilities for controllability in a broader sense, such as platform-side controllability or controllability with multiple objectives, which are not thoroughly defined and categorized. Additionally, some surveys on explainable information retrieval (IR) have expressed concerns related to controllability. However, there remains a significant distinction between explainability and controllability. First, explainability primarily focuses on making the embedding-based retrieval or recommendation models transparent and interpretable. In contrast, the notion of controllability discussed in this work emphasizes enabling retrieval or recommendation models to generate results under control, in accordance with given specific requirements. For instance, a recent survey on explainable IR [80, 81] paraphrases explainability as the ability to provide users or system designers with reasons clarifying why certain items are retrieved or recommended. Second, explainability can be regarded as a preliminary step toward achieving controllability. By offering such explanations, explainable IR facilitates human understanding of algorithmic decisions—whether the human is an end user or a system designer. This, in turn, enhances key aspects such as transparency, persuasiveness, effectiveness, and controllability. In summary, the key scientific challenge in the field of explainable IR is how to make the process and results of IR systems interpretable. Moreover, strong explainability can promote the research and development of CL.
@@
Additionally, these surveys are also relatively early, and many new controllable techniques (such as hypernetworks [82], large language models [73], etc.) have emerged since then. Considering the limitations of existing surveys, in this survey, we provide a formal definition of CL, and discuss CL and its applications in information retrieval.
1.5 Contribution
The current landscape of research and application deployment in the aforementioned areas motivates us to write a comprehensive survey on CL. As shown in Figure 1, this survey will encompass various aspects including the formal definition of CL, methods and applications in information retrieval, evaluations, challenges, and future directions.
  • Formal definition of CL. The survey will begin by providing a clear and concise formal definition of CL.
  • Applications of CL in IR. We will then delve into the applications of CL in the field of IR, exploring who controls, what is controllable, and how control is implemented in IR systems.
  • Future directions and challenges. Finally, the survey will discuss the future directions and potential challenges in the field of CL.
Overall, the survey on CL will serve as a valuable resource for students, researchers, practitioners, and policymakers interested in understanding the principles, applications, and future prospects of this important area of trustworthy machine learning.
2 Foundations of Controllable Learning (CL)